OpenStudy (laronjaxon):

please help What is the solution to the equation 7 –3x ≈ 9 ?

OpenStudy (anonymous):

You want to isolate x. First subtract 7 to both sides. What do you get?

OpenStudy (anonymous):

7-9=3x -2=3x x=-2/3

OpenStudy (anonymous):

Hey @manishsatywali Thanks for trying to help but giving answers away is against the code of conduct!

OpenStudy (laronjaxon):

A = 0.376 Bx = 0.295 C = –0.295 Dx = –0.376

OpenStudy (anonymous):

@RyanL. that guy is asking for solution

OpenStudy (anonymous):

@manishsatywali That doesn't make it ok you are not allowed to give answers at all.

OpenStudy (anonymous):

@manishsatywali See the person doesn't even know what answer choice to pick from when you clearly gave that person an answer.

OpenStudy (laronjaxon):

>.< its okay guys ! relax. Many people get answers/help I really could use it

OpenStudy (anonymous):

He gave you the answer. If you can't figure out what answer choice to pick from then you really need to study.

OpenStudy (anonymous):

@RyanL. i need coc instruction cn u message me them all?....or tell me wher i cn find them? coz i really nt understanding wat actually is there in coc